Please help me finish solving problem.
Solve by the substitution method.
4x+3y= -9
x=6-9y
------------
4(6-9y) +3y= -9
24-54y+3y= -34
Y=?
X=?

Answer by stanbon(75887)

Solve by the substitution method.
4x+3y= -9
x=6-9y
----------
Substitute x = 6-9y into the 1st equation to solve for "y":
4(6-9y) + 3y = -9
24 - 36y + 3y = -9
-33y = -33
y = 1
-------------------
Since x = 6-9y, x = 6 -9*1 = -3
--------------
Ans: (-3,1)
======================================================